OpenStudy (anonymous):

solve tan2x - square root of 3 = 0 for principal values of x, and express the solutions in degrees. Please be very specific in the steps you show :) The concept is really new to me, and I want to understand why you do what.

OpenStudy (anonymous):

\[\tan 2x=\sqrt{3}=\tan \frac{ \pi }{ 3 },2x=\frac{ \pi }{ 3 },x=?\]

OpenStudy (anonymous):

sorry, where does the tan pi/3 come from?

OpenStudy (anonymous):

\[\tan 2x=\sqrt{3}=\tan 60,2x=60,x=?\]

OpenStudy (anonymous):

\[\tan(whatever)=\sqrt3\\ whatever =\frac{\pi}{3}\]

OpenStudy (anonymous):

ohhh ok that makes so much more sense now!

OpenStudy (anonymous):

so then one of the solutions would be x= 30 degrees?

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Oh?

OpenStudy (anonymous):

well yes, but it really is \(\frac{\pi}{6}\)

OpenStudy (anonymous):

Ohh ok. Thank you. Are there other solutions too, or is it just that one?

OpenStudy (anonymous):

tangent is periodic, there are an infinite number of solutions